# Applying the maximum entropy principle to neural networks enhances multi-species distribution models

### v3

@@ D.1 Batch size effect on model performance
The mean AUC values for small batch sizes such as batch size 10 averaged 0.765. Increasing the batch size to 25 does not improve the overall AUC. Even though very small, the improvement comes mainly from AWT, NSW, and SWI. The average AUC in all regions shows better values for batch sizes of 100 and 250, with values of 0.767. Conversely, when the batch size increases to 1000 or 2500, the average AUC values for all regions decrease, reaching 0.766 and 0.764, respectively.
When analysing specific regions, we notice that regions like AWT and NSW show high variability in their AUC values. AWT exhibits the most instability, as its AUC value fluctuates between 0.702 and 0.716 across different batch sizes, suggesting that it may require a larger batch size for consistent performance. Conversely, SWI maintains a high AUC value around 0.848 with minimal changes, indicating that it is relatively unaffected by batch size variations. Similarly, regions such as CAN, NSW and SA also show limited impact, with AUC values remaining close to 0.732, 0.751 and 0.801, respectively, regardless of the batch size.
For some regions, such as SWI and CAN, the effects of batch size appear to be minimal, with good approximations obtained regardless of batch size. However, a medium batch size can improve computational efficiency on machines, providing a balance between performance and resource utilisation compared to larger batch sizes. On the other hand, the AWT remains the most unstable region, indicating that it requires more precise parametrisation to achieve optimal performance.
D.2 Influence of hidden layer number on model performance
@@
D.3 Sensitivity to the L2 regularisation weight
@@
Table 9: AUC values by bird species in Canada (Ontario) dataset according to different loss functions. The best average AUC for each species (row) is highlighted in bold, while the second-best AUC value is underlined.
spid count class Batch size
10 25 100 250 1000 2500
can01 16 rare 0.943 0.942 0.941 0.942 0.942 0.942
can02 740 abundant 0.709 0.708 0.707 0.707 0.705 0.706
can03 165 rare 0.658 0.655 0.655 0.659 0.662 0.670
can04 42 rare 0.755 0.754 0.741 0.750 0.767 0.786
can05 138 rare 0.656 0.657 0.653 0.651 0.647 0.638
can06 27 rare 0.886 0.886 0.886 0.886 0.886 0.885
can07 221 rare 0.695 0.692 0.690 0.687 0.673 0.663
can08 322 common 0.552 0.554 0.556 0.557 0.555 0.547
can09 119 rare 0.731 0.733 0.735 0.735 0.724 0.714
can10 234 rare 0.700 0.700 0.703 0.703 0.696 0.691
can11 478 common 0.594 0.591 0.590 0.592 0.590 0.590
can12 312 rare 0.848 0.847 0.846 0.846 0.846 0.846
can13 39 rare 0.701 0.695 0.687 0.682 0.681 0.689
can14 18 rare 0.923 0.921 0.918 0.919 0.920 0.921
can15 721 abundant 0.622 0.618 0.627 0.636 0.638 0.644
can16 57 rare 0.858 0.853 0.844 0.843 0.848 0.853
can17 313 rare 0.786 0.789 0.789 0.794 0.799 0.801
can18 612 common 0.762 0.760 0.764 0.765 0.779 0.792
can19 109 rare 0.482 0.488 0.494 0.511 0.519 0.498
can20 380 common 0.778 0.779 0.777 0.776 0.778 0.782
Table 10: AUC values by bird species in Canada (Ontario) dataset according to different weight decay values.
spid count class w=0 w=1e-6 w=3e-6 w=1e-5 w=3e-5 w=1e-4 w=3e-4 w=1e-3 w=3e-3 w=1e-2 w=3e-2 w=1e-1 w=3e-1
can01 16 rare 0.941 0.941 0.941 0.941 0.942 0.942 0.943 0.942 0.943 0.943 0.895 0.729 0.541
can02 740 abundant 0.706 0.706 0.705 0.705 0.705 0.706 0.706 0.708 0.716 0.721 0.717 0.702 0.688
can03 165 rare 0.652 0.652 0.653 0.653 0.654 0.656 0.661 0.673 0.674 0.667 0.662 0.627 0.579
can04 42 rare 0.728 0.728 0.731 0.731 0.734 0.743 0.764 0.801 0.783 0.765 0.710 0.547 0.544
can05 138 rare 0.649 0.650 0.649 0.649 0.650 0.649 0.648 0.639 0.625 0.621 0.611 0.554 0.520
can06 27 rare 0.887 0.887 0.887 0.887 0.886 0.886 0.886 0.886 0.884 0.883 0.845 0.736 0.666
can07 221 rare 0.683 0.683 0.682 0.682 0.682 0.682 0.678 0.663 0.657 0.662 0.662 0.621 0.602
can08 322 common 0.562 0.562 0.562 0.562 0.561 0.560 0.555 0.539 0.525 0.521 0.526 0.526 0.522
can09 119 rare 0.739 0.739 0.738 0.738 0.737 0.735 0.728 0.704 0.689 0.699 0.712 0.674 0.604
can10 234 rare 0.705 0.705 0.705 0.704 0.704 0.702 0.697 0.687 0.679 0.668 0.656 0.611 0.556
can11 478 common 0.590 0.590 0.590 0.590 0.591 0.591 0.591 0.592 0.595 0.595 0.592 0.575 0.558
can12 312 rare 0.844 0.844 0.844 0.844 0.845 0.846 0.846 0.848 0.850 0.848 0.844 0.829 0.775
can13 39 rare 0.651 0.652 0.652 0.652 0.656 0.665 0.682 0.712 0.734 0.732 0.690 0.621 0.581
can14 18 rare 0.902 0.902 0.903 0.904 0.907 0.913 0.919 0.922 0.923 0.921 0.772 0.582 0.555
can15 721 abundant 0.618 0.617 0.620 0.620 0.624 0.629 0.640 0.656 0.675 0.696 0.674 0.605 0.568
can16 57 rare 0.825 0.825 0.826 0.826 0.830 0.836 0.847 0.856 0.856 0.854 0.817 0.759 0.611
can17 313 rare 0.779 0.780 0.781 0.781 0.785 0.791 0.799 0.800 0.788 0.770 0.757 0.708 0.630
can18 612 common 0.760 0.760 0.761 0.761 0.762 0.764 0.775 0.795 0.799 0.798 0.789 0.759 0.690
can19 109 rare 0.534 0.535 0.535 0.536 0.536 0.533 0.521 0.474 0.461 0.469 0.481 0.483 0.487
can20 380 common 0.776 0.776 0.776 0.776 0.776 0.775 0.776 0.784 0.780 0.768 0.758 0.709 0.621
Table 11: Impact of hidden layer number on DeepMaxent performance with L2 regularisation (w=1e-2). The table presents average AUC and Pearson coefficient values across all regions.
@@
Table 12: Impact of batch size on DeepMaxent performance across all regions. The table presents average AUC and Pearson coefficient values across all regions.
Batch size avg AUC avg Pearson Coefficient
@@
Table 13: Comparison of method performance of DeepMaxent using TGB and using a equal weight between species by region-averaged AUC and averaged over all regions. The best average AUC for each column is highlighted in bold, while the second-best averaged AUC is underlined
@@
Table 14: Comparison of performance using DeepMaxent without TGB approach according to different weight decay values. The highest average AUC for each column is highlighted in bold, with the second-highest average AUC underlined.

### v4

@@ D.1 Sensitivity to the mini-batch size
The mean AUC values for small mini-batch sizes such as mini-batch size 10 averaged 0.765. Increasing the mini-batch size to 25 does not improve the overall AUC. Even though very small, the improvement comes mainly from AWT, NSW, and SWI. The average AUC in all regions shows better values for mini-batch sizes of 100 and 250, with values of 0.767. Conversely, when the mini-batch size increases to 1000 or 2500, the average AUC values for all regions decrease, reaching 0.766 and 0.764, respectively.
When analysing specific regions, we notice that regions like AWT and NSW show high variability in their AUC values. AWT exhibits the most instability, as its AUC value fluctuates between 0.702 and 0.716 across different mini-batch sizes, suggesting that it may require a larger mini-batch size for consistent performance. Conversely, SWI maintains a high AUC value around 0.848 with minimal changes, indicating that it is relatively unaffected by mini-batch size variations. Similarly, regions such as CAN, NSW and SA also show limited impact, with AUC values remaining close to 0.732, 0.751 and 0.801, respectively, regardless of the mini-batch size.
For some regions, such as SWI and CAN, the effects of mini-batch size appear to be minimal, with good approximations obtained regardless of mini-batch size. However, a medium mini-batch size can improve computational efficiency on machines, providing a balance between performance and resource utilisation compared to larger mini-batch sizes. On the other hand, the AWT remains the most unstable region, indicating that it requires more precise parametrisation to achieve optimal performance.
D.2 Sensitivity to the number of hidden layers
@@
D.3 Sensitivity to the number of neurons per hidden layer
Table 10 presents the impact of the number of neurons per hidden layer on the performance of DeepMaxent on the NCEAS dataset, all other hyper-parameters kept the same as for in main results, measured by AUC values across regions: AWT, CAN, NSW, NZ, SA, and SWI.
Hidden Size AUC across all region
50 0.75853
100 0.76583
250 0.76775
500 0.76662
750 0.76548
Table 10: AUC value as a function of hidden layer size for deepMaxent models.
D.4 Sensitivity to the L2 regularisation weight
@@
Table 12: Impact of hidden layer number on DeepMaxent performance. The table presents average AUC and Pearson coefficient values across all regions.
@@
Table 13: Impact of mini-batch size on DeepMaxent performance across all regions. The table presents average AUC and Pearson coefficient values across all regions.
Mini-batch size avg AUC avg Pearson Coefficient
@@
Table 14: Comparison of method performance of DeepMaxent using TGB and using a equal weight between species by region-averaged AUC and averaged over all regions.
@@
Table 15: Comparison of performance using DeepMaxent without TGB approach according to different weight decay values.
@@
Table 16: Impact of weight decay value on performances of baseline losses. The table presents average AUC and Pearson coefficient values across all regions.
Weight decay value loss AUC avg Pearson Coefficient
no BCE 0.763 0.233
3e5 BCE 0.764 0.235
1e4 BCE 0.764 0.239
3e4 BCE 0.763 0.240
1e3 BCE 0.755 0.231
3e3 BCE 0.751 0.208
no CE 0.744 0.162
3e5 CE 0.743 0.161
1e4 CE 0.744 0.162
3e4 CE 0.744 0.162
1e3 CE 0.744 0.162
3e3 CE 0.744 0.162
no Poisson loss 0.758 0.241
3e5 Poisson loss 0.759 0.242
1e4 Poisson loss 0.759 0.243
3e4 Poisson loss 0.757 0.241
1e3 Poisson loss 0.748 0.228
3e3 Poisson loss 0.721 0.198
D.5 Ablation study of MLP components in DeepMaxent
To better understand the contributions of architectural components in the DeepMaxent MLP model, we conducted an ablation study using the optimal hyper-parameters identified in the main experiments. We evaluated the model under four variants:
  • Full model: with skip connections and activation functions (ReLU).
  • No skip connections: activation functions kept.
  • No activation functions: skip connections kept.
  • No skip connections and no activation functions: a baseline MLP.
All models were retrained on the same dataset (NCEAS) with the same training procedure over ten same seeds. Performance was assessed via average AUC over species.
Table 17: Ablation study of MLP components in DeepMaxent across different regions. Models trained with optimal hyper-parameters.
Regions
Model variant AWT CAN NSW NZ SA SWI avg
Full model 0.712 0.732 0.752 0.753 0.806 0.850 0.768
No skip connections 0.714 0.731 0.752 0.747 0.802 0.848 0.766
No activations 0.704 0.729 0.738 0.736 0.763 0.831 0.750
No skip connections, no activations 0.702 0.731 0.738 0.741 0.764 0.830 0.750
The full model, incorporating both skip connections and activation functions, achieves the best average performance (0.768). Removing skip connections results in a slight marginal decrease, indicating that their contribution is beneficial but moderate. On the other hand, removing activation functions has a greater impact on performance, with a more pronounced decrease in the average AUC (0.750). These results highlight the importance of activation functions for the network’s learning ability, while skip connections play a secondary but positive role.
These results highlight the importance of activation functions for the learning capacity of the network, reflecting the non-linear nature of the data. In addition, the flexibility provided by residual connections (skip connections) improves model optimisation by facilitating gradient propagation and enabling more stable and efficient learning of complex representations, which contributes to better generalisation.
D.6 The importance of multi-species modelling in DeepMaxent: a case study in Switzerland
DeepMaxEnt was tested on each species independently over the SWI region, keeping the same set of parameters to build single-species models. In both cases, we applied the TGB correction to account for potential biases. For each species, we trained the model separately and evaluated its performance using the AUC metric. To compare these results with a multi-species approach, we calculated the average of the AUC values obtained for all the single-species models.
Table 18: Comparison of DeepMaxEnt performance between single-species and multi-species models in the Swiss region. The table reports AUC values for each approach, with results averaged across all species.
Method avg AUC for SWI
Single-species DeepMaxent 0.820
Multi-species DeepMaxent 0.850
Maxent (with TGB) 0.837
Table 18 shows the results for the single-species and multi-species models. It shows that the multi-species model performs much better. This may be explained by several reasons: (i) The set of hyper-parameters optimal for the multi-species model may not be suited for each single species model, and (ii) learning a feature extractor across multiple species is more robust one independent feature extractor for each individual species, as already showed in the past (e.g. Botella et al. [2018]). Adjusting hyper-parameters individually for each species could potentially improve performance of the single-species DeepMaxent models, but this approach would add up an important computational overload to the already important computational cost of fitting individual deep models.
Appendix E Statistical significance of performance differences
To assess whether the observed differences in predictive performance are statistically significant, we conducted pairwise Wilcoxon signed-rank tests between DeepMaxent and three alternative loss functions: Poisson, BCE and CE. These tests were applied to the macro AUC scores obtained over 10 random seeds, ensuring a fair comparison under identical experimental conditions.
As shown in Table 19, the resulting p-values are both below 0.002, indicating that the improvements achieved by DeepMaxent over both BCE and CE are statistically significant at the 0.01 level. The Wilcoxon statistic equals 0.0 in both cases, which further supports the consistent results of DeepMaxent across all tested seeds.
Table 19: Wilcoxon signed-rank tests over 10 random seeds comparing DeepMaxent to BCE and CE in terms of macro AUC. The low p-values indicate that the observed differences are statistically significant.
Comparison Wilcoxon Statistic p-value
DeepMaxent vs BCE 0.0 0.00195
DeepMaxent vs CE 0.0 0.0019
DeepMaxent vs Poisson 0.0 0.0019
Appendix F Comparison of performances across abundance classes on the NCEAS dataset
Figure 6: Comparison of average AUC values for AWT region by loss and abundance classes
Figure 7: Comparison of average AUC values for CAN region by loss and abundance classes
Figure 8: Comparison of average AUC values for NSW region by loss and abundance classes
Figure 9: Comparison of average AUC values for NZ region by loss and abundance classes
Figure 10: Comparison of average AUC values for SA region by loss and abundance classes
Figure 11: Comparison of average AUC values for SWI region by loss and abundance classes
[…]
Table 20: Mean value and standard deviation across all regions by species abundance class across all loss functions.
species class
Loss common intermediate rare
[…]
Appendix G Estimated probabilities maps by species with corresponding PA data, for CAN region
Figure 12: Estimated relative probabilities for the species can02 (a common species): (a) Presence-Only (PO) data, yellow points; (b) Presence-Absence (PA) data where green corresponds to presences, and magenta to absences. (c–f) Estimated from different loss functions: (c) DeepMaxent, (d) Binary Cross-Entropy (BCE), (e) Cross-Entropy (CE), (f) Poisson loss.
Presence/Absence Estimated relative probability
can01
Refer to caption Refer to caption
can02
Refer to caption Refer to caption
can03
Refer to caption Refer to caption
can04
Refer to caption Refer to caption
can05
Refer to caption Refer to caption
Figure 13: Locations of species presences (green) and absences (magenta) based on PA observations, alongside predicted values from the reference DeepMaxent model across species can01 to can05
Presence/Absence Estimated relative probability
can06
Refer to caption Refer to caption
can07
Refer to caption Refer to caption
can08
Refer to caption Refer to caption
can09
Refer to caption Refer to caption
can10
Refer to caption Refer to caption
Figure 14: Locations of species presences (green) and absences (magenta) based on PA observations, alongside predicted values from the reference DeepMaxent model across species can06 to can10
Presence/Absence Estimated relative probability
can11
Refer to caption Refer to caption
can12
Refer to caption Refer to caption
can13
Refer to caption Refer to caption
can14
Refer to caption Refer to caption
can15
Refer to caption Refer to caption
Figure 15: Locations of species presences (green) and absences (magenta) based on PA observations, alongside predicted values from the reference DeepMaxent model across species can11 to can15
Presence/Absence Estimated relative probability
can16
Refer to caption Refer to caption
can17
Refer to caption Refer to caption
can18
Refer to caption Refer to caption
can19
Refer to caption Refer to caption
can20
Refer to caption Refer to caption
Figure 16: Locations of species presences (green) and absences (magenta) based on PA observations, alongside predicted values from the reference DeepMaxent model across species can16 to can20
Mean Std
Reference
Refer to caption Refer to caption
Batch size
Smaller value
Refer to caption Refer to caption
Batch size
Higher value
Refer to caption Refer to caption
Weight decay
Smaller value
Refer to caption Refer to caption
Weight decay
Higher value
Refer to caption Refer to caption
Figure 17: Estimated mean and standard deviation of cumulative estimated probabilities across all species in the CAN region calculated over 10 random seeds under different hyper-parameter settings (batch size and weight decay).
Appendix H Computational time
Table 21 summarizes the computational time of the different training procedures on the GeoPlant dataset (>4 thousand species, 5 million occurrences) on a laptop. The results indicate that the GPU implementation significantly accelerates the learning of DeepMaxent, whereas the CPU-based training required a considerably longer execution times. Additionally, even though Maxent was fitted with only 10,000 sites drawn among the millions of background sites, training DeepMaxent on a CPU (including all background sites) was twice quicker.
Method Time
Maxent (CPU) $\sim$ 6.9 hours
DeepMaxent (CPU) $\sim$ 3 hours
DeepMaxent (GPU) $\sim$ 40 minutes
Table 21: Approximate training time for each model on the GeoPlant dataset. MaxEnt is trained as a single-species model, whereas DeepMaxent is trained jointly on all species for 100 epoch on CPU and GPU
Appendix I Dataset description
I.1 NCEAS
Figure 18: Location of presence-only (PO) occurrences for all 20 bird species in Canada
I.2 GeoPlant
Figure 22: A map representing 1% of GeoPlant PO observations.
Figure 23: A map representing 1% of GeoPlant PA data.
Figure 24: A cube representation of Landsat time series for a single observation at a specific location (latitude, longitude). The cube has dimensions of 6 spectral bands × 4 seasons × 21 years, where each value represents the seasonal median of a given spectral band for a specific year